HUBBLE SPACE TELESCOPE DAILY REPORT #2218 PERIOD COVERED: 0000Z (UTC) 09/28/98 - 0000Z (UTC) 09/29/98 Daily Status Report as of 272/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of STIS/CCD/MA2/WF/PC-2/NIC/1/2/3 8058 (HDF-South Observations) The Space Telescope Imaging Spectrograph (CCD and MA2), the WF/PC-2, and the Near Infrared Camera (Cameras One, Two and Three) were all used to make long exposure Hubble Deep Field (HDFS-22333) observations. This program is designed to provide deep observations of a high-galactic latitude field in the southern continuous viewing zone. The STIS field is centered on a bright QSO at z=2.25. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 4 Successful : 4 2.3 Operations Notes: The two-week-long Hubble Deep Field-South (HDF-S) campaign began with today's observations. These observations are described above. No substantial problems were encountered. The NSSC-1 status buffer was dumped and reset three times per ROP NS-3. These status buffer events were expected as a result of HDF-S observations. They will occur at each SAA exit until the STIS event flag 2 is reset at the end of the HDF-S campaign. Using ROP SR-1, the SSR EDAC error counter was reset four times. As documented in HSTAR 6695, the STIS MAMA-2 MCE reset at 271/194053Z during an SAA passage and while the MCE was in transition to low voltage. The STIS flight software error counter was reset per ROP NS-12. The MAMA-2 was recovered via the SMS at the time (271/2205735Z) of the next high voltage on commanding. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH